please help me solve this problem please thank you
The ancient Greeks though that the most pleasing shape for a rectangle was one for which the ratio of length to width was 8 to 5, the golden ratio. If the length of a rectangular painting is 2 ft longer than its width, then for what dimensions would the length and width have the golden ratio?

Answer by stanbon(75887)

The ancient Greeks though that the most pleasing shape for a rectangle was one for which the ratio of length to width was 8 to 5, the golden ratio. If the length of a rectangular painting is 2 ft longer than its width, then for what dimensions would the length and width have the golden ratio?
---------
length/width = 8/5
---------------------
(x+2)/x = 8/5
5(x+2) = 8x
8x-5x = 10
3x = 10
x = 3 1/3 ft (width)
x+2 = 5 1/3 ft (length)
=============================
Cheers,
Stan H.